The continuous compounding rate is 13,974% per year. The equivalent nominal rate, compounded quarterly, is


1
Expert's answer
2022-03-08T00:03:02-0500

Solution


Let us consider, the equivalent nominal rate is ""x"", then the future value is


"A = P{\left( {1 + \frac{x}{n}} \right)^{nt}}\"


But for the continuous compounding, we have


"A = P \cdot {e^{rt}}"


Therefore,

"P \cdot {e^{rt}} = P{\left( {1 + \frac{r}{n}} \right)^{nt}}\\"


"{\left( {{e^r}} \right)^t} = {\left( {1 + \frac{r}{n}} \right)^{nt}}\\"


"{e^r} = {\left( {1 + \frac{r}{n}} \right)^n}\\"


"{e^{0.13974}} = {\left( {1 + \frac{x}{4}} \right)^4}\\"


"1 + \frac{x}{4} = {\left( {{e^{0.13974}}} \right)^{{\textstyle{1 \over 4}}}}\\"


"\frac{x}{4} = {\left( {{e^{0.13974}}} \right)^{{\textstyle{1 \over 4}}}} - 1\\"


"x = 4 \cdot \left[ {{{\left( {{e^{0.13974}}} \right)}^{{\textstyle{1 \over 4}}}} - 1} \right]\\"


"x = 0.1422095828"


Hence the nominal rate is "14.221 \%"